# Sparse Ellipsoidal Radial Basis Function Network for Point Cloud Surface Representation

### v2

@@ 3.2 Ellipsoid RBF Neural Network
A function of the form $\Phi(\bm{x})=\Phi(\|\bm{x}-\bm{c}\|)$ is called a RBF, its value depends only on the radius from $\bm{x}$ to the center $\bm{c}$. There are many RBF expressions, one example is given by:
$\Phi(\bm{x})=e^{-\|\bm{x}-\bm{c}\|^{2}},$ (1)
@@
In this work, we extend the classical RBF by adopting a more flexible ellipsoidal form. The ERBF is defined as follows:
@@
$\bm{R}\left(\theta_{x}\right)=\left[\begin{array}[]{ccc}0&\cos\theta_{x}&-\sin\theta_{x}\\ 0&\sin\theta_{x}&\cos\theta_{x}\end{array}\right],$
@@
$\Psi(\bm{x}_{i},\bm{\widetilde{\Theta}})$ $=\sum_{j=1}^{M}w_{j}^{2}\hat{\Phi}(\bm{x}_{i},\Theta_{j})$ (4)
$=\sum_{j=1}^{M}w_{j}^{2}e^{-\left\|\bm{D}_{j}\bm{R}_{j}\left(\bm{x}_{i}-\bm{c}_{j}\right)\right\|_{2}^{2}},$
@@
Suppose that the set of grid points $\bm{X}$ used to perform the point cloud surface reconstruction consists of $N$ points, that is, $\bm{X}=\displaystyle\bigcup_{i=1}^{N}\bm{x}_{i}$, where $\bm{x}_{i}$ denotes the $i$-th point. Next, according to Eq. (5), each point $\bm{x}_{i}$ is fed into SE-RBFNet to obtain the corresponding predicted SDF value, the output of SE-RBFNet is denoted as $\bm{O}$, as follows:
@@
For convenience of notation, matrices $\bm{F}$ are introduced in Eq. (5). $\bm{F}$, denoted as the feature matrix, has dimensions $N\times M$, and each element represents the response value of the $i$-th point coordinate $\bm{x}_{i}$ under the influence of the $j$-th ERBF $\hat{\Phi}(\bm{x}_{i},\Theta_{j})$. Subsequently, the marching cubes algorithm [23] is applied to the output $\bm{O}$ to extract the surface.
@@ 3.3 Loss Function Design
Given an input point cloud $P=\left\{\bm{p}_{i}\right\}_{i=1}^{K_{p}}$, $K_{p}$ is the number of points in the point cloud, we first build an adaptive octree structure to hierarchically partition the 3D space. The octree recursively subdivides each cell into eight smaller cubes until a prespecified resolution is reached or the minimum number of points per cell is satisfied. The coordinates of each cubic vertex serve as grid points, and the set of grid point coordinates is denoted by $G=\left\{\bm{g}_{j}\right\}_{j=1}^{K_{g}}$, $K_{g}$ is the number of grid points. The signed distance $S(\bm{g}_{j})$ for each grid point $\bm{g}_{j}$ is calculated from the point cloud $P$. A variety of methods have been developed to compute the SDF, which mainly differ in how the distance and the sign are determined. Classical approaches compute the SDF as the distance to the closest surface point with signs determined by normal orientation [57] or ray casting [24]. Advanced techniques, such as the generalized winding number [58], improve robustness in complex geometries, while learning-based methods including DeepSDF [3], SAL [59], and Neural-Pull [51] directly predict SDFs from point clouds. In addition, several open-source libraries, such as Kaolin [60] and point-cloud-utils [61], provide built-in functions for efficient SDF computation. The surface itself is characterized by the zero-level set of the signed distance: $\{\bm{g}_{j}\in\mathbb{R}^{3},S(\bm{g}_{j})=0\}$. For points in the inside (outside) of the surface, $S(\bm{g}_{j})$ is negative (positive). As can be seen from Eq. (4), the output of $\Psi$ is nonnegative. Therefore, to fit the signed distance using the SE-RBFNet, we scale all signed distances $S$ to [0, 3] using the following nonlinear transformation.
$\hat{S}$ $=3\cdot e^{-h(S-m)^{2}}\textrm{,}$ (6)
where
$m$ $=min(S)\textrm{,}$
$h$ $=\frac{\log 3}{m}.$
Eq. (6) normalizes $S$ values less than 0 (inside the surface) to the interval [1, 3] and values greater than 0 (outside the surface) to the interval [0, 1]. Notably, the zero-level set of $S=0$ (i.e., the reconstructed surface) is mapped to $\hat{S}=1$.
For the network model, we merge the grid points $G$ and the point cloud coordinates $P$ as the total training points, denoted as $V$, $V=\{G,P\}=\left\{\bm{v}_{i}\right\}_{i=1}^{N}$, $N=K_{p}+K_{g}$ is the total number of training points, $\bm{v}_{i}$ denotes the $i$-th training point. $\hat{S}$ is merged with the all-1 vector $I$ as the label of $P$, denoted as $T$, $T=\{\hat{S},I\}=\left\{t_{i}\right\}_{i=1}^{N}$, $t_{i}$ is the $i$-th SDF value. Here the length of $I$ is equal to the number of point cloud points. The SE-RBFNet model are used to appoximate the given SDF values of $V$ as follows:
@@
Based on Eq. (9), Eq. (8) can be simplified as follows: $\mathcal{L}=\alpha\mathcal{L}_{2}+\beta\mathcal{L}_{1}$, where $\mathcal{L}_{2}$ ensures reconstruction accuracy of the surface by minimizing the squared error between predicted and ground-truth values on training points $V$. $\mathcal{L}_{1}$ enforces sparsity in the network parameters by penalizing the $\mathcal{L}_{1}$ norm of the weights, encouraging a sparse solution for $\bm{W}$. Weights in $\bm{W}$ that are close to zero correspond to basis elements with negligible contributions to the network, which can safely be removed. $\alpha>0$ and $\beta>0$ are dynamic weighting coefficients that balance the two loss terms. Our aim is to minimize the loss function $\mathcal{L}$ and find the corresponding optimal network parameters, $\bm{\widetilde{\Theta}}$.
However, in surface reconstruction, more attention is paid to the points close to the surface, where $\left\{\bm{x}\in\mathbb{R}^{3}\,\middle|\,\Psi(\bm{x})=1\right\}$. To further reduce the computational cost, two thresholds, $\tau_{1}$ (with $\tau_{1}<1$) and $\tau_{2}$ (with $\tau_{2}>1$), each close to 1, are introduced to filter the training points for loss computation at each iteration, as defined by the following formula:
@@
In Eq. (10), in addition to sampling training points near the surface, we also sample points where the predicted SDF significantly deviates from the ground truth. Specifically, $id_{0}$ selects from both interior and exterior regions close to the surface, which are crucial for accurately approximating the SDF near the surface and ensuring effective surface representation through ERBFs. $id_{1}$ ($id_{2}$) selects points with erroneous predictions in the exterior (interior) region to help the network learn and optimize further, avoiding artificial bulges or dents on the surface represented by the predicted SDF. For instance, $id_{1}$ ($id_{2}$) prevents misclassification of points that should properly reside outside (inside) the surface as being erroneously close to or crossing the surface boundary. Only the points corresponding to $id_{all}$ are subject to loss calculation, so Eq. (8) can be rewritten as:
@@
During the optimisation of the loss function, the $\mathcal{L}_{1}$ regularization term is not applied at every iteration. There are two prerequisites that must be met. One is that, starting from epoch $e_{p}\geq k^{e}$, the standard deviation of the $\mathcal{L}_{2}$ loss over the last $k^{e}$ epochs satisfies the convergence condition:
@@
where $\operatorname{std}(\cdot)$ denotes standard deviation operation, $\mathcal{L}_{2}^{(e)}$ represents the $\mathcal{L}_{2}$ loss at epoch $e$, $e_{p}$ indicates the current training epoch. The parameters $k^{l_{2}}$ and $\tau^{l_{2}}$ denote the number of iterations for $\mathcal{L}_{2}$ loss convergence judgement and the corresponding threshold, respectively. Once this condition of Eq. (12) is satisfied, the $\mathcal{L}_{1}$ is considered to be active. However, whether sparse optimization is actually applied further depends on a second condition:
$\max_{i\in id_{all}}\left(\Psi(\bm{v}_{i},\bm{\widetilde{\Theta}})-t_{i}\right)^{2}<\tau_{m},$ (13)
where $\tau_{m}$ is a hyperparameter for comparison with the maximum squared error. If either of these conditions is not satisfied, the optimization proceeds using only the $\mathcal{L}_{2}$ loss i.e., $\alpha=1\text{},\beta=0$.
When performing sparse optimisation with the $\mathcal{L}_{1}$ regularization term, we adopt an adaptive weighting mechanism [21] instead of fixed coefficients. This mechanism is inspired by the Frank-Wolfe method for Pareto multi-task learning and dynamically adjusts the trade-off between the two objectives—accuracy and sparsity—during the optimization process. This dynamic approach is formulated as:
@@
where $[\cdot]_{+,{}_{\tau}^{1}}$ represents clipping to $[0,1]$ as $[\sigma]_{+,{}_{\tau}^{1}}=\max(\min(\sigma,1),0)$. This dynamic multi-objective optimization strategy automatically adjusts the trade-off between accuracy and sparsity throughout training (step 20 in Algorithm 3), leading to improved generalization and robustness.
@@ 3.4.1 Inscribed Sphere-Based Initialization for ERBF
The hidden layer of SE-RBFNet consists of multiple ellipsoidal radial basis functions. To effectively initialize the network parameters of SE-RBFNet, we use the largest inscribed sphere method [62] to calculate the initialization parameters of each ERBF. This method ensures that the initialization process can adapt to the geometry of the target surface, facilitating the learning of the SDF used for surface reconstruction.
For interior points $G_{\text{in}}=\left\{g_{i}\mid 1<\hat{S}_{i}<3\right\}$ and the corresponding signed distance $\hat{S}_{\text{in}}$, we iteratively detect the maximal inscribed spheres using Algorithm 1 to obtain the initial centers $C=\left\{\bm{c}_{j}\right\}_{j=1}^{M}$ and the initial weights $W=\left\{w_{j}\right\}_{j=1}^{M}$. Meanwhile, in order to visually demonstrate the calculation process of the inscribed sphere, Figure1 (d) illustrates the inner and outer points (left), as well as the distribution of the maximum inscribed spheres (right). As shown, the inscribed spheres densely cover the internal space of the point cloud, providing an effective approximation of its geometric structure.
From Algorithm 1, the weights and centers are initialized. Then, the shape parameters $\bm{D}_{j},j=1,2,\ldots,M$ are initialized with identical diagonal elements, i.e., $d_{j1}=d_{j2}=d_{j3}$. Assume that the Gaussian distribution is satisfied between each Gaussian kernel center and its nearest neighboring Gaussian kernel center. To minimize the mutual influence between Gaussian kernels, we ensure that the function value of each Gaussian kernel at the position of its nearest neighboring kernel is sufficiently small. This effectively reduces the interference between different kernels as well. Under this assumption, $d_{j1}$ is computed as follows:
$w_{j}^{2}e^{-d_{j1}^{2}\left(\hat{d}_{i}/2\right)^{2}}=\gamma.$ (15)
Here, $\gamma$ is a threshold value chosen to minimize the mutual influence between kernels. $\hat{d}_{j}=\min\left\|c_{j}-c_{i}\right\|,i=1,2,\ldots,M$ and $i\neq j$. This means that $\hat{d}_{j}$ is the shortest distance from the current kernel center $\bm{c}_{j}$ to any other kernel center. After simplification, the $d_{j1}$ can be derived from Eq. (15) as:
$d_{j1}=d_{j2}=d_{j3}=\sqrt{\frac{-\ln\left(\frac{\gamma}{w_{j}^{2}}\right)}{\left(\hat{d}_{j}/2\right)^{2}}}.$ (16)
Initialise $\bm{D}_{j}=\operatorname{diag}\left(d_{j1},d_{j2},d_{j3}\right)$. Initialize the rotation angle of each ellipsoid along the principal axis to 0, the corresponding of each $\bm{R}_{j}$ is the identity matrix.
Algorithm 1 Maximum Inscribed Spheres Computation
Input: interior gird points $G_{\text{in}}$, the corresponding signed distance $\hat{S}_{\text{in}}$, point cloud $P$.
Compute the closest distance of each interior point to the point cloud surface, denoted as $\bm{d}^{in}$.
@@ 3.4.2 Hierarchical Optimization
In this section, we describle a hierarchical optimization method for training the aforementioned SE-RBFNet on the octree grid points. Specifically, we perform the training process on the octree grid points layer by layer. Assuming that $G^{i}$ represents the grid point of the $i$-th layer of the octree, $G$ can be represented as follows:
@@
and
$G^{i}$ =$=\{G^{i}_{1}\textrm{,}G^{i}_{2}\textrm{,}...\textrm{,}G^{i}_{k_{i}}\},$
where ${k_{i}}$ indicates the number of grid points in the $i$-th layer, and the total number of layers in the octree is $l$. Similarly, the signed distance corresponding to each grid point is also layered, that is,
@@
and
$\hat{S}^{i}$ =$=\{\hat{S}^{i}_{1}\textrm{,}\hat{S}^{i}_{2}\textrm{,}...\textrm{,}\hat{S}^{i}_{k_{i}}\}.$
During optimization, the training process begins with the grid points of the coarse layer, and data from finer layers is progressively incorporated. The specific steps are as follows:
@@
    Combine $\hat{G}^{l_{s}}$ with the point cloud points $P$ to form $V^{l_{s}}=\{\hat{G}^{l_{s}}\textrm{,}P\}$ as the initial training set, send it to the model along with the corresponding label $T^{l_{s}}=\{\hat{S}^{1}\textrm{,}\hat{S}^{2}\textrm{,}...\textrm{,}\hat{S}^{l_{s}}\textrm{,}I\}$ for supervised optimization.
@@
    Calculate the loss using Eq. (8) and continue optimizing until convergence. Then add the points of the next fine layer to the training set.
@@
Note that the formulae presented here and the formulas below are based on the initial layer $l_{s}$, but the same strategy is applied when optimizing subsequent layers. In practice, the initial layer $l_{s}$ is typically set to the third-to-last layer.

### v4

@@ 3.2 Ellipsoid RBF Neural Network
A function of the form $\Phi(\bm{x,c})=\Phi(\|\bm{x}-\bm{c}\|)$ is called an RBF, whose value depends only on the distance from $\bm{x}$ to the center $\bm{c}$. There are many RBF expressions. One example is given by:
$\Phi(\bm{x,c})=e^{-\|\bm{x}-\bm{c}\|^{2}},$ (1)
@@
In this work, the ERBF [63] extends the classical RBF by adopting a more flexible ellipsoidal form. The ERBF in $\mathbb{R}^{3}$ is defined as follows:
@@
$\bm{R}\left(\theta_{x}\right)=\left[\begin{array}[]{ccc}1&0&0\\ 0&\cos\theta_{x}&-\sin\theta_{x}\\ 0&\sin\theta_{x}&\cos\theta_{x}\end{array}\right],$
@@
$\Psi(\bm{x}_{i},\bm{\widetilde{\Theta}})=\sum_{j=1}^{M}w_{j}^{2}\hat{\Phi}(\bm{x}_{i},\Theta_{j})=\sum_{j=1}^{M}w_{j}^{2}e^{-\left\|\bm{D}_{j}\bm{R}_{j}\left(\bm{x}_{i}-\bm{c}_{j}\right)\right\|_{2}^{2}},$ (4)
@@
Suppose the set of grid points used for surface extraction is denoted by $\bm{X}$, consisting of $N$ points: $\bm{X}=\displaystyle\bigcup_{i=1}^{N}\bm{x}_{i}$, where $\bm{x}_{i}$ denotes the $i$-th point. Next, according to Eq. (5), each point $\bm{x}_{i}$ is fed into SE-RBFNet to obtain the corresponding predicted SDF value. The output of SE-RBFNet is denoted as $\bm{O}$, as follows:
@@
For convenience of notation, we introduce the feature matrix $\bm{F}\in\mathbb{R}^{N\times M}$ in Eq. (5), where each element of $\bm{F}$ corresponds to the response value of the $i$-th point $\bm{x}_{i}$ under the influence of the $j$-th ERBF $\hat{\Phi}(\bm{x}_{i},\Theta_{j})$. Subsequently, the marching cubes algorithm [34] is applied to the output $\bm{O}$ to extract the explicit surface.
@@ 3.3 Loss Function Design
Given a set of sampling points on an implicit surface $P=\left\{\bm{p}_{i}\right\}_{i=1}^{K_{p}}$, $K_{p}$ is the number of points, we assume that the corresponding octree grid points $G=\left\{\bm{g}_{j}\right\}_{j=1}^{K_{g}}$ and the signed distance value $S(\bm{g}_{j})$ at each grid point are provided as input, where $K_{g}$ is the number of grid points. The surface itself is characterized by the zero-level set of the signed distance: $\{\bm{g}_{j}\in\mathbb{R}^{3},S(\bm{g}_{j})=0\}$. For points inside (outside) the surface, $S(\bm{g}_{j})$ is negative (positive). As can be seen from Eq. (4), the output of $\Psi$ is nonnegative. Therefore, to fit the signed distance using the SE-RBFNet, we scale all signed distances $S$ to [0, 2] using the following nonlinear transformation.
$\hat{S}$ $=2\cdot e^{-h(S-m)^{2}}\textrm{,}$ (6)
where $m=min(S)$, $h=\frac{\ln(2)}{m^{2}}$. Eq. (6) normalizes $S$ values less than 0 (inside the surface) to the interval [1, 2] and values greater than 0 (outside the surface) to the interval [0, 1]. Notably, the zero-level set of $S=0$ (i.e., the reconstructed surface) is mapped to $\hat{S}=1$.
For the network model, we merge the grid points $G$ and the sampling points $P$ as the total training points, denoted as $V$, $V=\{G,P\}=\left\{\bm{v}_{i}\right\}_{i=1}^{N}$, $N=K_{p}+K_{g}$ is the total number of training points, $\bm{v}_{i}$ denotes the $i$-th training point. $\hat{S}$ is merged with the all-1 vector $I$ as the label of $P$, denoted as $T$, $T=\{\hat{S},I\}=\left\{t_{i}\right\}_{i=1}^{N}$, $t_{i}$ is the $i$-th SDF value. Here, the length of $I$ is equal to the number of sampling points. The SE-RBFNet model is used to approximate the given SDF values of $V$ as follows:
@@
Based on Eq. (9), Eq. (8) can be rewritten concisely as follows: $\mathcal{L}=\alpha\mathcal{L}_{2}+\beta\mathcal{L}_{1}$, where $\mathcal{L}_{2}$ ensures approximate accuracy by minimizing the squared error between predicted and ground-truth values on training points $V$. $\mathcal{L}_{1}$ enforces sparsity in the network parameters by penalizing the $\mathcal{L}_{1}$ norm of the weights, encouraging a sparse solution for $\bm{W}$. Weights in $\bm{W}$ that are close to zero correspond to basis elements with negligible contributions to the network, which can safely be removed. $\alpha\geq 0$ and $\beta\geq 0$ are dynamic weighting coefficients that balance the two loss terms. Our aim is to minimize the loss function $\mathcal{L}$ and find the corresponding optimal network parameters, $\bm{\widetilde{\Theta}}$.
However, in surface representation, more attention is paid to the points close to the surface, where $\left\{\bm{x}\in\mathbb{R}^{3}\,\middle|\,\Psi(\bm{x})=1\right\}$. To further reduce the computational cost, two thresholds, $\tau_{1}$ (with $\tau_{1}<1$) and $\tau_{2}$ (with $\tau_{2}>1$), each close to 1, are introduced to filter the training points for loss computation at each iteration, as defined by the following formula:
@@
In Eq. (10), $id_{0}$ selects from both interior and exterior regions close to the surface, which are crucial for accurately approximating the SDF near the surface and ensuring effective surface representation through ERBFs. $id_{1}$ ($id_{2}$) selects points with erroneous predictions in the exterior (interior) region to help the network learn and optimize further, avoiding artificial bulges or dents on the surface represented by the predicted SDF. For instance, $id_{1}$ ($id_{2}$) prevents misclassification of points that should properly reside outside (inside) the surface as being erroneously close to or crossing the surface boundary. Only the points corresponding to $id_{all}$ are subject to loss calculation, so Eq. (8) can be rewritten as:
@@
During the optimization of the loss function, the $\mathcal{L}_{1}$ regularization term is applied only when two conditions are simultaneously satisfied. First, starting from epoch $e_{p}\geq k^{l_{2}}$, the $\mathcal{L}_{2}$ loss is considered converged, meaning that its standard deviation over the last $k^{l_{2}}$ epochs falls below a threshold $\tau^{l_{2}}$:
@@
where $\operatorname{std}(\cdot)$ denotes the standard deviation operation, $\mathcal{L}_{2}^{(e)}$ represents the $\mathcal{L}_{2}$ loss at epoch $e$, $e_{p}$ indicates the current training epoch. Second, the maximum absolute error between the predicted and the ground-truth SDFs must be below a threshold $\tau_{m}$:
$\max_{i\in id_{all}}\left|\Psi(\bm{v}_{i},\bm{\widetilde{\Theta}})-t_{i}\right|<\tau_{m},$ (13)
If either of these conditions is not satisfied, the optimization proceeds using only the $\mathcal{L}_{2}$ loss, i.e., $\alpha=1\text{},\beta=0$.
When performing sparse optimization with the $\mathcal{L}_{1}$ regularization term, we adopt a dynamic weighting mechanism [50] instead of using fixed coefficients $\alpha$ and $\beta$ in Eq. (11). This mechanism is inspired by the Frank-Wolfe method for Pareto multi-task learning [50] and dynamically adjusts the trade-off between accuracy and sparsity during the optimization process. Specifically, at each optimization step, the coefficient $\alpha$ is computed based on the gradients of the two loss terms, reflecting their current relative contributions, and $\beta$ is set as $1-\alpha$. This dynamic approach is formulated as:
@@
where $\cdot]_{+,{}_{\tau}^{1}}$ represents clipping to $[0,1]$ as $\sigma]_{+,{}_{\tau}^{1}}=\max(\min(\sigma,1),0)$. This dynamic multi-objective optimization strategy (step 19 in Algorithm 3) automatically adjusts the coefficients during training, balancing accuracy and sparsity according to the model’s state, which improves generalization and robustness.
@@ 3.4.1 Inscribed Sphere-Based Initialization for ERBF
The hidden layer of SE-RBFNet consists of multiple ERBFs. To effectively initialize the network parameters of SE-RBFNet, we use the largest inscribed sphere method [9] to calculate the initialization parameters of each ERBF. This method ensures that the initialization process can adapt to the geometry of the target surface, facilitating the learning of the SDF used for surface representation.
For interior points $G_{\text{in}}=\left\{g_{i}\mid 1<\hat{S}_{i}<2\right\}$ and their corresponding signed distance $\hat{S}_{\text{in}}$, we iteratively detect the maximal inscribed spheres using Algorithm 1 to obtain the initial centers $C=\left\{\bm{c}_{j}\right\}_{j=1}^{M}$ and the initial weights $W=\left\{w_{j}\right\}_{j=1}^{M}$. To visually demonstrate the calculation process of the inscribed sphere, Figure 1.(d) illustrates the inner and outer points (left), as well as the distribution of the maximum inscribed spheres (right). As depicted, the inscribed spheres densely cover the internal space of the implicit surface, providing an effective approximation of its geometric structure.
Algorithm 1 Inscribed Sphere-Based Initialization for ERBF
Input: interior gird points $G_{\text{in}}$, the corresponding signed distance $\hat{S}_{\text{in}}$, sampling points $P$.
Compute the closest distance of each interior point to $P$, denoted as $\bm{d}^{in}$.
@@
From Algorithm 1, the weights and centers are initialized. Then, the shape parameters $\bm{D}_{j},j=1,2,\ldots,M$ are initialized with identical diagonal elements, i.e., $d_{j1}=d_{j2}=d_{j3}$. Assume that the Gaussian distribution is satisfied between each Gaussian kernel center and its nearest neighboring Gaussian kernel center. To minimize the mutual influence between Gaussian kernels, we ensure that the function value of each Gaussian kernel at the position of its nearest neighboring kernel is sufficiently small. This effectively reduces the interference between different kernels as well. Under this assumption, $d_{j1}$ satisfies the following equation:
$w_{j}^{2}e^{-d_{j1}^{2}\left(\hat{d}_{i}/2\right)^{2}}=\gamma.$ (15)
Here, $\gamma$ is a threshold value chosen to minimize the mutual influence between kernels. $\hat{d}_{i}=\min\left\|c_{j}-c_{i}\right\|_{2},i=1,2,\ldots,M$ and $i\neq j$. This means that $\hat{d}_{i}$ is the shortest distance from the current kernel center $\bm{c}_{j}$ to any other kernel center. After simplification, the $d_{j1}$ can be derived from Eq. (15) as:
$d_{j1}=d_{j2}=d_{j3}=\frac{2\sqrt{-\ln\left(\gamma/w_{j}^{2}\right)}}{\hat{d}_{i}}.$ (16)
Initialize $\bm{D}_{j}=\operatorname{diag}\left(d_{j1},d_{j2},d_{j3}\right)$. The rotation angle of each ellipsoid along the principal axis is set to zero, so that each corresponding rotation matrix $\bm{R}_{j}$ is initialized as the identity matrix.
@@ 3.4.2 Hierarchical Optimization
In this section, we describe a hierarchical optimization method for training the aforementioned SE-RBFNet on multi-level grid points. Although the method is applicable to any hierarchical grid structure, we use an octree as an example for illustration. Specifically, we perform the training process on the octree grid points layer by layer. Assuming that $G^{i}$ represents the grid point of the $i$-th layer of the octree, $G$ can be represented as follows:
@@
and $G^{i}=\{G^{i}_{1}\textrm{,}G^{i}_{2}\textrm{,}...\textrm{,}G^{i}_{k_{i}}\}$, where $k_{i}$} indicates the number of grid points in the $i$-th layer, and the total number of layers in the octree is $l$. Similarly, the signed distance corresponding to each grid point is also layered, that is,
@@
and $\hat{S}^{i}=\{\hat{S}^{i}_{1}\textrm{,}\hat{S}^{i}_{2}\textrm{,}...\textrm{,}\hat{S}^{i}_{k_{i}}\}$. During optimization, the training process begins with the grid points of the coarse layer, and data from finer layers is progressively incorporated. The specific steps are as follows:
@@
    Combine $\hat{G}^{l_{s}}$ with the sampling points $P$ to form $V^{l_{s}}=\{\hat{G}^{l_{s}}\textrm{,}P\}$ as the initial training set, send it to the model along with the corresponding label $T^{l_{s}}=\{\hat{S}^{1}\textrm{,}\hat{S}^{2}\textrm{,}...\textrm{,}\hat{S}^{l_{s}}\textrm{,}I\}$ for supervised optimization.
@@
    Using Eq. (11), compute the loss and optimize until convergence, then add the points of the next finer layer to the training set.
@@
Note that the formulations presented here, as well as those introduced below, are derived with respect to the initial layer $l_{s}$, while the same optimization strategy is applied to subsequent layers. In practice, the initial layer $l_{s}$ is typically set to the third-to-last layer.